OpenStudy (mtwizzle420):

What is the length of the hypotenuse of the right triangle shown. Triangle will be in comments below. @jordanloveangel

OpenStudy (mtwizzle420):



OpenStudy (anonymous):

are you sure your triangle is correct?I can see the side length is greater than the hypotenuse

OpenStudy (anonymous):

oh my goodness, i had to do a review on this a few weeks ago...its so frustrating ;(

OpenStudy (anonymous):

is their an actual picture or no cause you can attach file

OpenStudy (mtwizzle420):

The 10 and c are switched my bad!!! My computer is being janky guys im so sorry!

OpenStudy (anonymous):

A^2+B^2+C ^2 so 10^2+31^=what @mtwizzle420

OpenStudy (anonymous):

10^2+31^2=

OpenStudy (mtwizzle420):

100+961=1061

OpenStudy (anonymous):

10^2+31^2= 1061 Take the square root of 1061 and you will get C

OpenStudy (mtwizzle420):

530.5 is what I got?

OpenStudy (mtwizzle420):

530.5?

OpenStudy (zale101):